OpenStudy (anonymous):

Find delta such that |square root of (5x +1)-4>0.5 if 0<|x-3|

OpenStudy (blockcolder):

\[|x-3|<\delta \Rightarrow |\sqrt{5x+1}-4|<0.5\] \[\sqrt{5x+1}-4\cdot\frac{\sqrt{5x+1}+4}{\sqrt{5x+1}+4}=\frac{5x+1-16}{\sqrt{5x+1}+4}=\frac{5x-15}{\sqrt{5x+1}+4}=\frac{5(x-3)}{\sqrt{5x+1}+4}\] Use this to help you find suitable bounds for delta.